OpenStudy (firejay5):

Solve the equation for the specified variable. Please be clear on and instructions: E = MC^2; solve for m

OpenStudy (anonymous):

Okey, what we want is to isolate M. The equation should look something like this: M = .... If we divide both sides with C^2 we get: E / C^2 = MC^2 / C^2 The C^2 in the numerator and the denominator will cancel leaving us with E / C^2 = M Or M = E / C^2

OpenStudy (firejay5):

This is the answer: E / C^2 = M Or M = E / C^2

OpenStudy (anonymous):

Well either one, the are the same. But just pick one. The second is best in my opinion. M = E / C^2

OpenStudy (firejay5):

thank you best answer I gave you a medal

OpenStudy (anonymous):

Thank you :)

OpenStudy (firejay5):

you are welcome dude